A 3000-kg car pulled a 1000-kg crate along a straight highway. A driving force of 9000 N accelerates the car and the crate at 1.5 m/s2 . Assuming that one-third of the total friction is acting on the crate and two-thirds on the car, determine the total friction and the magnitude of the tension at the rope


1
Expert's answer
2021-06-07T15:01:48-0400
"(m+M)a=F-F_f\\(3000+1000)1.5=9000-F_f\\F_f=3000\ N"

"1000(1.5)=T-1000\\T=2500\ N"